# Identity-Preserving Text-to-Video Generation by Frequency Decomposition

### v2

@@ Abstract
Identity-preserving text-to-video (IPT2V) generation aims to create high-fidelity videos with consistent human identity. It is an important task in video generation but remains an open problem for generative models. This paper pushes the technical frontier of IPT2V in two directions that have not been resolved in the literature: (1) A tuning-free pipeline without tedious case-by-case finetuning, and (2) A frequency-aware heuristic identity-preserving Diffusion Transformer (DiT)-based control scheme. To achieve these goals, we propose ConsisID, a tuning-free DiT-based controllable IPT2V model to keep human-identity consistent in the generated video. Inspired by prior findings in frequency analysis of vision/diffusion transformers, it employs identity-control signals in the frequency domain, where facial features can be decomposed into low-frequency global features (e.g., profile, proportions) and high-frequency intrinsic features (e.g., identity markers that remain unaffected by pose changes). First, from a low-frequency perspective, we introduce a global facial extractor, which encodes the reference image and facial key points into a latent space, generating features enriched with low-frequency information. These features are then integrated into the shallow layers of the network to alleviate training challenges associated with DiT. Second, from a high-frequency perspective, we design a local facial extractor to capture high-frequency details and inject them into the transformer blocks, enhancing the model’s ability to preserve fine-grained features. To leverage the frequency information for identity preservation, we propose a hierarchical training strategy, transforming a vanilla pre-trained video generation model into an IPT2V model. Extensive experiments demonstrate that our frequency-aware heuristic scheme provides an optimal control solution for DiT-based models. Thanks to this scheme, our ConsisID achieves excellent results in generating high-quality, identity-preserving videos, making strides towards more effective IPT2V.
Figure 2: Overview of the proposed ConsisID. Based on Findings of DiT, low-frequency facial information is embedded into the shallow layers, while high-frequency information is incorporated into the vision tokens within the attention blocks. The ID-preserving Recipe is applied to ease training and improve generalization. The cross face, DropToken and Dropout are executed based on probability.
@@ 1 Introduction
Additionally, the above efforts are predominantly based on U-Net and cannot be adapted to the emerging DiT-based video model [26, 72, 62, 73, 60]. This challenge may stem from the inherent limitations of DiT compared to U-Net, including greater difficulty in training convergence and weakness in perceiving facial details. From some prior findings in frequency analysis of vision/diffusion transformers [42, 2, 3, 4, 53, 48, 67], we can know that the reason is: Finding 1: Shallow (e.g., low-level, low-frequency) features are essential for pixel-level prediction tasks in diffusion models, as they ease model training. U-Net facilitates model convergence by aggregating shallow features to the decoder via long skip connections, a mechanism that DiT does not incorporate; Finding 2: Transformers have limited perception of high-frequency information, which is important for preserving facial features. The encoder-decoder architecture of U-Net naturally possesses multi-scale features (e.g., richness in high-frequency), while DiT lacks a comparable structure. To develop a DiT-based control model, it is necessary to address these issues first.
@@
  • Extensive experiments demonstrate our ConsisID can generate high-quality, editable, consistent identity-preserving videos, benefiting from our frequency-aware identity-preserving T2V DiT-based control scheme.
@@ 2 Related Work
Tuning-based Identity-preserving T2V Models.    Diffusion models are widely recognized for their strong generative capabilities [19, 40, 41, 46, 37, 68, 69], significantly advancing the development of identity-preserving generative models [57, 8, 70, 34]. Initially, the researchers used tuning-based methods to generate content that matched the input ID. This process requires finetuning pretrained model for each new person during inference. For example, DreamBooth [44] introduced a novel loss function to fine-tune the entire network, embedding identity information while preserving the original generative capabilities. LoRA [21], similar to DreamBooth [44], requires training only a small subset of network parameters. In contrast, Textual Inversion [11] freezes the pretrained network and embeds identity information into a trainable word embedding. Subsequent tuning-based methods, including both image and video models based on U-Net or DiT architectures [45, 25, 58, 56, 59, 7], generally follow three main approaches. While these models demonstrate substantial effectiveness, the requirement to fine-tune for each new identity restricts their practical applicability.
@@ 3.1 Preliminaries
Diffusion Transformer.    The DiT-based video generation model shows significant potential in simulating the physical world [6, 62, 73]. Despite being a novel architecture, research on controllable generation has been limited, and current methods [71, 13, 9, 38] largely resemble U-Net based approaches [70, 34, 11]. However, no study has yet examined why this approach works with DiT. Drawing from prior analyses of Diffusion and Transformer from a frequency domain perspective [42, 2, 3, 4, 53, 48], we conclude that: (1) Low-level (e.g., shallow-layer) features are essential for pixel-level prediction tasks in diffusion models, which helps facilitate model training; (2) Transformers have limited perception for high-frequency information, which is important for controllable generation. Based on these, we decouple ID features into high- and low-frequency parts and inject them into specific locations, achieving effective identity-preserving text-to-video generation.
@@ 3.2.1 Low-frequency View: Global Facial Extractor
In light of Finding 1, enhancing low-level (e.g., shallow, low-frequency) features accelerates model convergence. To easily adapt a pre-trained model for the IPT2V task, the most direct approach is concatenating the reference face with the noise input latent [5]. However, the reference face contains both high-frequency details (e.g., eye and lip textures) and low-frequency information (e.g., facial proportions and contours). From Finding 2, prematurely injecting high-frequency information into the Transformer is inefficient and may hinder the model’s processing of low-frequency information, as the Transformer focuses primarily on low-frequency features. In addition, feeding the reference face directly into the model could introduce irrelevant noise such as lighting and shadows. To mitigate this, we extract facial key points, convert them to an RGB image, and then concatenate it with the reference image, as shown in Figure 2. This strategy focuses the model’s attention on the low-frequency signals in the face, while minimizing the impact of extraneous features. We found that when this component is discarded, the model has a gradient explosion. The objective function is changed to:
@@
where $\psi_{\theta}(\cdot)$ is the global facial extractor, $f$ represents the reference image.
@@ 3.2.2 High-frequency View: Local Facial Extractor
In light of Finding 2, we recognize that Transformers have limited sensitivity to high-frequency information. It can be concluded that relying solely on global facial features is insufficient for IPT2V generation, as global facial features primarily consist of low-frequency information and lack the intrinsic features necessary for editing. This task requires not only maintaining identity consistency, but also incorporating editing capabilities, such as generating videos of faces with the same identity but varying age and makeup. Achieving this requires the extraction of facial features that are unaffected by non-ID attributes (e.g., expression, posture, and shape), since age and makeup do not alter a person’s core identity. We define these features as intrinsic identity features (e.g., high-frequency).
Previous research [16, 15, 14] use local features from the CLIP image encoder [39] as intrinsic features to improve editing capabilities. However, since CLIP is not specifically trained on face datasets, the extracted features contain harmful non-face information [29, 54, 63]. Therefore, we choose to use a face recognition backbone [10] to extract intrinsic identity features. Instead of using the output of the backbone as the intrinsic identity feature, we use the penultimate layer, which retains more spatial information related to identity. However, these features still lack sufficient semantic information [13, 14], which is crucial for personalized video generation.
To address these issues, we first use a facial recognition backbone to extract features that are strong in the intrinsic identity representation, and a CLIP image encoder to capture features that are strong in semantics. We then use the Q-former [27, 28, 61] to fuse these two features, producing intrinsic identity features enriched with high-frequency semantic information. To reduce the impact of irrelevant features from CLIP, dropout [1, 22] is applied before entering into Q-Former. Additionally, we concatenate the shallow, multi-scale features from the facial recognition backbone, after interpolation, with the CLIP features. This method ensures that the model effectively captures essential intrinsic identity features while filtering out external noise unrelated to identity. After extracting the intrinsic id features, we apply cross-attention to interact with the visual tokens produced by each attention block of the pre-trained model, effectively enhancing the high-frequency information in DiT:
@@ 4.1 Setup
Implementation details.    ConsisID selects DiT-based generation architectures CogVideoX-5B [62] as our baseline for validation. We use an in-house human-centric dataset for training, which differs from previous datasets [55, 64, 35] that focus only on the face. In the training phase, we set the resolution to 480$\times$720 and extracted 49 consecutive frames at a stride of 3 from each video as training data. We set the batch size to 80, the learning rate to $3\times 10^{-6}$, and the total number of training steps to 1.8k. The classify free guidance random null text ratio is set to $0.1$, with AdamW serving as an optimizer and cosine_with_restarts as a learning rate scheduler. The training strategy is the same as Section 3.2.3. We set $\alpha$ and $\beta$ in the dynamic cross-face loss ($\mathcal{L}_{e}$) and dynamic mask loss ($\mathcal{L}_{f}$) to $0.5$, respectively. In the inference phase, we employ DPM [49] with a sampling step of $50$, and a text-guidance ratio of $6.0$.
@@
Benchmark.    Since there is an absence of an evaluation dataset, we select 30 persons who were not included in the training data and sourced five high-quality images for each ID from the internet. We then design 90 distinct prompts, encompassing a variety of expressions, actions, and backgrounds for evaluation. Building on previous works [38, 15], we evaluate four dimensions: (1). Identity Preservation: We use FaceSim-Arc [10] and introduce FaceSim-Cur, which assesses identity preservation by measuring feature differences between face regions in the generated videos and those in real face images within the ArcFace [10] and CurricularFace [23] feature spaces. (2) Visual Quality: We utilize FID [18] by calculating feature differences in the face regions between the generated frames and real face images within the InceptionV3 [51] feature space. (3) Text Relevance: We utilize CLIPScore [17] to measure the similarity between the generated videos and the input prompts. (4). Motion Amplitude: Due to the lack of reliable metrics, we evaluate through the user study.
@@ 4.3 Quantitative Analysis
We present a comprehensive quantitative evaluation of different methods, with results displayed in Table 1. Consistent with Figure 5, our method outperforms state-of-the-art methods across five metrics. For identity preservation, ConsisID achieves a higher score by designing appropriate identity signals for DiT from a frequency perspective. By contrast, ID-Animator [15] is not optimized for IPT2V and only partially retains facial features, resulting in lower FaceSim-Arc [10] and FaceSim-Cur scores. For Text Relevance, ConsisID not only controls expressions via prompts but also adjusts actions and backgrounds, achieving higher CLIPScore [17]. Regarding visual quality, the FID is presented solely as a reference due to its limited alignment [69, 31, 36, 30] with human perception. Please refer to Figure 5 and 4 for qualitative analysis of the visual quality.
@@ 4.5 Effect of the Identity Signal Injection in DiT
To assess the effectiveness of Finding 1 and Finding 2, we perform ablation experiments on different methods of injecting control signals into DiT. Specifically, these experiments involved (a) injecting only low-frequency face information with key points into the noise latent, (b) injecting only high-frequency face signals within the attention block, (c) combining (a) and (b), (d) based on (c), but the low-frequency face information does not contain key points, and (e - f) based on (c), but the high-frequency signal is injected at the output or input of the attention block. (g) injecting only high-frequency face signals before the attention block. The results are shown in Figure 7 and Table 3. For Finding 1, we observe that only injecting high-frequency signals (a) greatly increases the training difficulty, causing the model to fail to converge due to the lack of low-frequency signal injection. In addition, the inclusion of facial key points (d) allows a greater focus on low-frequency information, thereby facilitating training and improving model performance. For Finding 2, when only low-frequency signals are injected (b), the model lacks high-frequency information. This reliance on low-frequency signals causes the generated face in the video to copy the reference image, making it difficult to control facial expressions, movements, and other features through prompts. Furthermore, injecting identity signals into the attention block input (f - g) disrupts the intended frequency domain distribution of DiT, resulting in a gradient explosion. Embedding control signals in the attention block (c) is preferable to embedding them in the output (e) because attention block processes predominantly low-frequency information. By embedding high-frequency information internally, the attention block is guided to highlight intrinsic facial features, whereas injecting it into the output merely concatenates features without directing focus, reducing DiT’s modeling capacity. Moreover, we apply a Fourier transform to the generated videos (only the face region) to visually compare the influence of different components to extract facial information. As shown in Figure 3, the Fourier spectrum and the log amplitude of the Fourier transform reveal that injecting high or low-frequency signals can indeed enhance the corresponding frequency information of the generated face. Moreover, the low-frequency signal can be further enhanced by matching with the face key points, and injecting the high-frequency signal into the attention block has the highest feature utilization rate. Our method (c) shows strongest high and low frequency, further validating the efficiency benefit from Findings 1 and 2. To reduce overhead, for each identity, we only select 2 reference images for the evaluation.
@@
Figure 8: Effect of the Inversion Steps $t$. Overall quality does not improve consistently as $t$ increases, but first improves and then declines. This may be because the early steps are dominated by low frequency, whereas the later steps are dominated by high frequency.

### v3

@@ Abstract
Identity-preserving text-to-video (IPT2V) generation aims to create high-fidelity videos with consistent human identity. It is an important task in video generation but remains an open problem for generative models. This paper pushes the technical frontier of IPT2V in two directions that have not been resolved in the literature: (1) A tuning-free pipeline without tedious case-by-case finetuning, and (2) A frequency-aware heuristic identity-preserving Diffusion Transformer (DiT)-based control scheme. To achieve these goals, we propose ConsisID, a tuning-free DiT-based controllable IPT2V model to keep human-identity consistent in the generated video. Inspired by prior findings in frequency analysis of vision/diffusion transformers, it employs identity-control signals base on frequency domain, since facial features can be decomposed into low-frequency global features (e.g., profile, proportions) and high-frequency intrinsic features (e.g., identity markers that remain unaffected by pose changes). Extensive experiments demonstrate that our frequency-aware heuristic scheme provides an optimal control solution for DiT-based models, making strides toward more effective IPT2V.
Figure 2: Overview of the proposed ConsisID. Based on Findings of DiT, low-frequency facial information is embedded into the shallow layers, while high-frequency information is incorporated into the vision tokens within the attention blocks. The ID-preserving Recipe is applied to ease training and improve generalization. The cross face, DropToken, and Dropout are executed based on probability.
@@ 1 Introduction
Additionally, the above efforts are predominantly based on U-Net and cannot be adapted to the emerging DiT-based video model [37, 96, 83, 97, 76]. This challenge may stem from the inherent limitations of DiT compared to U-Net, including greater difficulty in training convergence and weakness in perceiving facial details. From some prior findings in frequency analysis of vision/diffusion transformers [53, 3, 4, 5, 66, 61, 88], we can know that the reason is: Finding 1: Shallow (e.g., low-level, low-frequency) features are essential for pixel-level prediction tasks in diffusion models, as they ease model training. U-Net facilitates model convergence by aggregating shallow features to the decoder via long skip connections, a mechanism that DiT does not incorporate; Finding 2: Transformers have limited perception of high-frequency information, which is important for preserving facial features. The encoder-decoder architecture of U-Net naturally possesses multi-scale features (e.g., richness in high-frequency), while DiT lacks a comparable structure. To develop a DiT-based control model, these must be addressed first. Please see Appendix for more details.
@@
  • Extensive experiments demonstrate that our ConsisID can generate high-quality, editable, consistent identity-preserving videos, benefiting from our frequency-aware identity-preserving T2V DiT-based control scheme.
@@ 2 Related Work
Tuning-based Identity-preserving T2V Models.    Diffusion models are widely recognized for their strong generative capabilities [23, 52, 43, 44, 45, 42, 89, 90], significantly advancing the development of identity-preserving generative models [71, 11, 91, 47]. Initially, the researchers used tuning-based methods to generate content that matched the input ID. This process requires finetuning pretrained model for each new person during inference. For example, DreamBooth [57] introduced a novel loss function to fine-tune the entire network, embedding identity information while preserving the original generative capabilities. LoRA [25], similar to DreamBooth [57], requires training only a small subset of network parameters. In contrast, Textual Inversion [15] freezes the pretrained network and embeds identity information into a trainable word embedding. Subsequent tuning-based methods, including both image and video models based on U-Net or DiT architectures [58, 33, 72, 70, 73, 10, 75, 82, 95], generally follow three main approaches. While these models demonstrate substantial effectiveness, the requirement to fine-tune for each new identity restricts their practical applicability.
@@ 3.1 Preliminaries
Diffusion Transformer.    The DiT-based video generation model shows significant potential in simulating the physical world [8, 83, 97]. Despite being a novel architecture, research on controllable generation has been limited, and current methods [94, 17, 12, 50] largely resemble U-Net based approaches [91, 47, 15]. However, no study has yet examined why this approach works with DiT. Drawing from prior analyses of Diffusion and Transformer from a frequency domain perspective [53, 3, 4, 5, 66, 61], we conclude that: (1) Low-frequency (e.g., shallow-layer) features are essential for pixel-level prediction tasks in diffusion models, which helps facilitate model training; (2) Transformers have limited perception for high-frequency information, which is important for controllable generation. Based on these, we decouple ID features into high- and low-frequency parts and inject them into specific locations, achieving effective identity-preserving text-to-video generation.
@@ 3.2.1 Low-frequency View: Global Facial Extractor
In light of Finding 1, enhancing low-level (e.g., shallow, low-frequency) features accelerates model convergence. To easily adapt a pre-trained model for the IPT2V task, the most direct approach is concatenating the reference face with the noise input latent [7]. However, the reference face contains both high-frequency details (e.g., eye and lip textures) and low-frequency information (e.g., facial proportions and contours). From Finding 2, prematurely injecting high-frequency information into the Transformer is inefficient and may hinder the model’s processing of low-frequency information, as the Transformer focuses primarily on low-frequency features. In addition, feeding the reference face directly into the model could introduce irrelevant noise such as lighting and shadows. To mitigate this, we extract facial key points, convert them to an RGB image, and then concatenate it with the reference image, as shown in Figure 2. This strategy focuses the model’s attention on the low-frequency signals in the face, while minimizing the impact of extraneous features. We found that when this component is discarded, the model is difficult to convergen. The objective function is changed to:
@@
where $\psi_{\theta}(\cdot)$ is a variational autoencoder, $f$ represents the reference image, we ignore key points here for simplicity.
@@ 3.2.2 High-frequency View: Local Facial Extractor
In light of Finding 2, we recognize that Transformers have limited sensitivity to high-frequency information. It can be concluded that relying solely on global facial extractor is insufficient for IPT2V generation, as global facial features lack of high-frequency information. So we use a face recognition backbone [13] to extract high-frequency features, as these are invariant to non-ID attributes (e.g., expression, posture, and shape). We refer to these features as intrinsic identity features (e.g., high-frequency), since age and makeup do not alter an individual’s core identity. Following [18], we utilize the penultimate layer of the backbone, rather than its output, as it retains more spatial information pertinent to identity. However, our experiments reveal that while the face recognition backbone improves identity consistency, it lacks the semantic features required for editing. This task demands not only maintaining identity consistency but also incorporating the ability to edit, such as generating videos of faces with the same identity but varying age and makeup. Previous research [20, 19] relies solely on the CLIP encoder [51] to enable editing capabilities. However, since CLIP is not specifically trained on face datasets, the features it extracts include irrelevant non-face information, which can compromise identity fidelity [36, 68, 84].
To address these challenges, we first use a facial recognition backbone to extract features that strongly represent intrinsic identity, and a CLIP image encoder to capture semantically rich features. We then employ the Q-Former [34, 35, 77] to fuse these features, resulting in intrinsic identity representations enriched with high-frequency semantic information. To mitigate the impact of irrelevant features from CLIP, dropout [2, 26] is applied post-processing. Additionally, we follow [18] to concatenate the shallow, multi-scale features from the facial recognition backbone (after interpolation) with the CLIP features. This approach ensures that the model captures essential intrinsic identity features while filtering out extraneous noise unrelated to identity. Finally, we apply cross-attention to facilitate interaction between this feature set and the visual tokens produced by each attention block of the pre-trained model, thereby enhancing the high-frequency information in the DiT:
@@ 4.1 Setup
Implementation details.    ConsisID selects DiT-based generation architectures CogVideoX-5B [83] as our baseline for validation. We use an in-house human-centric dataset for training, which differs from previous datasets [69, 85, 48] that focus only on the face. In the training phase, we set the resolution to 480$\times$720 and extracted 49 consecutive frames at a stride of 3 from each video as training data. We set the batch size to 80, the learning rate to $3\times 10^{-6}$, and the total number of training steps to 1.8k. The randomly discarded text rate is set to $0.1$, with AdamW serving as an optimizer and cosine_with_restarts as a learning rate scheduler. The training strategy is the same as Section 3.2.3. We set $\alpha$ and $\beta$ in the dynamic cross-face loss ($\mathcal{L}_{e}$) and dynamic mask loss ($\mathcal{L}_{f}$) to $0.5$, respectively. In the inference phase, we employ DPM [62] with a sampling step of $50$, and a classifier free guidance ratio of $6.0$. For more detials and results, please refer to Appendix.
@@
Benchmark.    Since there is an absence of an evaluation dataset, we select 30 persons who were not included in the training data and sourced five high-quality images for each ID from the internet. We then design 90 distinct prompts, encompassing a variety of expressions, actions, and backgrounds for evaluation. Building on previous works [50, 19], we evaluate four dimensions: (1). Identity Preservation: We use FaceSim-Arc [13] and introduce FaceSim-Cur, which assesses identity preservation by measuring feature differences between face regions in the generated videos and those in real face images within the ArcFace [13] and CurricularFace [27] feature spaces. (2) Visual Quality: We utilize FID [22] by calculating feature differences in the face regions between the generated frames and real face images within the InceptionV3 [64] feature space. (3) Text Relevance: We utilize CLIPScore [21] to measure the similarity between the generated videos and the input prompts. (4). Motion Amplitude: Due to the lack of reliable metrics [29, 90], we evaluate through the user study.
@@ 4.3 Quantitative Analysis
We present a comprehensive quantitative evaluation of different methods, with results displayed in Table 1. Consistent with Figure 4, our method outperforms state-of-the-art methods across five metrics. For identity preservation, ConsisID achieves a higher score by designing appropriate identity signals for DiT from a frequency perspective. By contrast, ID-Animator [19] is not optimized for IPT2V and only partially retains facial features, resulting in lower FaceSim-Arc [13] and FaceSim-Cur scores. For Text Relevance, ConsisID not only controls expressions via prompts but also adjusts actions and backgrounds, achieving higher CLIPScore [21]. Regarding visual quality, the FID is presented solely as a reference due to its limited alignment [90, 40, 49, 39] with human perception. Please refer to Figure 4 and 3 for qualitative analysis of the visual quality.
@@ 4.5 Effect of the Identity Signal Injection in DiT
To assess the effectiveness of Finding 1 and Finding 2, we perform ablation studies on different methods of injecting control signals into DiT. Specifically, these experiments involved (a) injecting only low-frequency face information with key points into the noise latent, (b) injecting only high-frequency face signals within the attention block, (c) combining (a) and (b), (d) based on (c), but the low-frequency face information does not contain key points, and (e - f) based on (c), but the high-frequency signal is injected at the output or input of the attention block. (g) injecting only high-frequency face signals before the attention block. To reduce overhead, for each identity, we only select 2 reference images each with 90 text prompts for the evaluation. The results are shown in Figure 6 and Table 3. For Finding 1, we observe that only injecting high-frequency signals (a) greatly increases the training difficulty, causing the model to fail to converge due to the lack of low-frequency signal injection. In addition, the inclusion of facial key points (d) allows a greater focus on low-frequency information, thereby facilitating training and improving model performance. For Finding 2, when only low-frequency signals are injected (b), the model lacks high-frequency information. This reliance on low-frequency signals causes the generated face in the video to copy the reference image, making it difficult to control facial expressions, movements, and other features through prompts. Furthermore, injecting identity signals into the attention block input (f - g) disrupts the intended frequency domain distribution of DiT, resulting in a gradient explosion. Embedding control signals in the attention block (c) is preferable to embedding them in the output (e) because attention block processes predominantly low-frequency information. By embedding high-frequency information internally, the attention block is guided to highlight intrinsic facial features, whereas injecting it into the output merely concatenates features without directing focus, reducing DiT’s modeling capacity. Moreover, we apply a Fourier transform to the generated videos (only the face region) to visually compare the influence of different components to extract facial information. As shown in Figure 7, the Fourier spectrum and the log amplitude of the Fourier transform reveal that injecting high or low-frequency signals can indeed enhance the corresponding frequency information of the generated face. Moreover, the low-frequency signal can be further enhanced by matching with the face key points, and injecting the high-frequency signal into the attention block has the highest feature utilization rate. Our method (c) shows strongest high and low frequency, further validating the efficiency benefit from Findings 1 and 2.